# Evidence-based multimodal fusion on structured EHRs and free-text notes for ICU outcome prediction

### v1

@@ RESULTS
Regarding predictive evaluation metrics, the models achieved comparable BACC across the three fusion settings for both tasks. Notably, the model based on modalities outperformed those based on data types and data sources in F1, AUROC, and AUPRC metrics. In terms of reliability evaluation metrics, the model based on modalities demonstrated superior reliability across different backbone models.
@@
Figure 3: The evaluation of our framework on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
DISCUSSION
Principle findings
@@ DISCUSSION
In this section, we present our findings across four key areas: the advantages of using free-text notes for mortality and PLOS prediction, the effectiveness of evidence-based multimodal framework, the impact of various fusion settings, and the influence of encoder selection.
Benefits of free-text notes
@@
From Table 2, it is evident that integrating free-text notes with structured data enhances ICU outcome prediction, boosting predictive performance and reliability. Free-text notes provide information not included in structured EHR data, such as nursing details, physician documentation, and radiology reports after ICU admission. This suggests that free-text EHR notes and structured inputs can complement each other in predictive modeling, leading to improved performance.
Effectiveness of evidence-based multimodal framework
Our evidence-based framework demonstrated improved predictive performance and higher reliability compared to existing multimodal approaches. This improvement is likely due to the heterogeneous nature of EHRs, which often contain imprecise information from diverse sources. By effectively extracting and integrating informative evidence through modalities, our framework addresses these challenges.
@@
Additionally, the current multimodal approach achieved higher recall in both tasks (refer to Appendix E). High recall reflects the model’s ability to capture the most true positive instances, often at the cost of accepting more false positives. In ICU settings, this could result in the unnecessary utilization of medical resources and equipment. In contrast, our framework strikes a better balance by demonstrating higher precision and specificity, reducing false positives. This capability is critical to appropriately allocate medical resources to ICU patients in real need.
It is also noteworthy that the improvement in BACC (1.05% for mortality prediction and 1.02% for PLOS prediction) is smaller than the improvement in F1 score (9.74% for mortality prediction and 6.04% for PLOS prediction). The F1 score, which emphasizes the balance between precision and recall, is particularly valuable for assessing improvements in identifying positive samples within imbalanced datasets, especially when positive outcomes are rare. Given that ICU outcomes are a minority in our dataset, this suggests that our framework is particularly effective at correctly identifying these critical outcomes.
Impacts of different fusion settings
To explore the performance of our framework across different fusion settings, it is evident that models based on modalities achieved higher F1 scores but lower BACCs compared to the other two fusion settings. This discrepancy arises from the metrics’ focus: the F1 score emphasizes performance on the positive class by balancing precision and recall, while balanced accuracy (BACC) provides an overall assessment of recall across all classes. Thus, models based on modalities are more effective at identifying ICU outcomes, likely due to the enhanced integration of information from independent sources facilitated by evidence theory.
This observation aligns with our analysis of data independence. Figures 4 and 5 visualize the independence of structured features and four types of free-text notes, respectively. In Figure 4, correlation coefficients confirm that features within structured data are not independent. Meanwhile, Figure 5 shows that Radiology notes form a distinct cluster, while the other types of notes overlap, indicating a lack of independence among them. Additionally, our analysis highlights a general relationship between predictive performance and reliability.
Influence of encoders selection
The evaluations presented in Tables 2 and 3 reveal that the choice of encoders plays a critical role in determining predictive performance. Among the models assessed, the FT-Transformer stands out as highly effective for tabular data, aligning with the findings of 23. This effectiveness can be attributed to the transformer’s capability to capture complex relationships among transformed numerical and categorical features, which enhances its predictive power.
CONCLUSION
In this paper, we address the challenge of predicting ICU outcomes by introducing an evidence-based multimodal framework that models both structured EHR data and free-text EHR notes. Our framework transforms deep features extracted from these two modalities into evidence through evidence mapping module, which is then fused using Dempster’s rule to make final predictions. Through experiments on the MIMIC-III dataset, we demonstrate the effectiveness of our framework in managing heterogeneous multimodal EHR data, reducing false positives and potentially improving the allocation of medical resources in the ICU.
While this paper focuses on binary classification tasks, many clinical applications require solutions for multiclass tasks (e.g., disease diagnosis) and continuous regression tasks (e.g., survival prediction). These are equally important and relevant for advancing clinical practice. In the future, we plan to expand our framework by incorporating additional data modalities, such as time series and medical images, to provide deeper clinical insights and enhance model performance. We also aim to extend the framework to handle multimodal EHR multiclass tasks, offering valuable predictive guidance for complex clinical scenarios. Additionally, we intend to investigate regression tasks, leveraging the recently introduced Epistemic Random Fuzzy Set (ERFS) theory 32, 33 and further building on developments in evidential regression 34.
ACKNOWLEDGMENTS
APPENDIX
A  Dempster-Shafer (DS) Theory
Let $\Omega=\{\omega_{1},\omega_{2},\cdots,\omega_{M}\}$ be a finite set of hypotheses about some question, called the frame of discernment. Evidence about a variable taking values in $\Omega$ can be represented by a mass function: $2^{\Omega}$ to [0,1] such that $m(\emptyset)=0$ and
$\sum_{A\subseteq\Omega}m(A)=1.$ (1)
For any hypothesis $A\subseteq\Omega$, the quantity $m(A)$ is interpreted as a share of a unit mass of belief allocated to the hypothesis that the truth is in $A$, and which cannot be allocated to any strict subset of $A$ based on the available evidence. Set $A$ is called a focal set of $m$ if $m(A)>0$. A mass function is said to be Bayesian if its focal sets are singletons, and logical if it has only one focal set. Two mass functions $m_{1}$ and $m_{2}$ representing independent items of evidence can be combined conjunctively by Dempster’s combination rule 17 $\oplus$ as
for all $A\neq\emptyset$, where $\sum_{B\cap C=\emptyset}m_{1}(B)m_{2}(C)$ is the degree of conflict among the two pieces of evidence, The nice information fusion attribute of DS theory points out the high potential of using DS theory for heterogenetic medical data analysis.
After aggregating all available evidence, we adopt the pignistic transformation, as proposed by Smets in the Transferable Belief Model 35, to make a final decision based on the combined mass function. This framework bases decisions on the pignistic probability distribution with the following expression:
B  Evidence mapping (EM)
As illustrated in Figure 6, the evidence mapping module consists of one input layer, one hidden layer, and one output layer. The input layer is composed of $H$ units ($H$ is the number of prototypes), whose weights vectors are prototypes $\pi_{1},\pi_{2},\cdots,\pi_{H}$ in input space. The activation of unit $h$ in the input layer is
$s_{h}=\beta_{h}\mathrm{exp}(-\gamma_{h}d_{h}^{2}),$ (4)
Figure 6: The illustration of evidence mapping
$m_{h}(\{\omega_{c}\})$ $=u_{h}^{(c)}s_{h},\quad c=1,2,\cdots,M,$ (5a)
$m_{h}(\Omega)$ $=1-s_{h},$ (5b)
$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{H}\in\mathbb{R}^{M+1}.$ (7)
The combined mass functions (the outputs of the EM module) represent the degrees of belief about the given class with $m(\{\omega_{c}\})$, as well as its prediction uncertainty with $m(\Omega)$. In our binary classification case, the dimension of EM module outputs would be three.
C  Evidence fusion
$m=m_{1}\oplus m_{2}\oplus\cdots\oplus m_{K}\in\mathbb{R}^{M+1},$ (8)
D  The details of augmented learning algorithm
$\mathcal{L}_{p}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{c=1}^{M}w_{c}y_{i,c}\mathrm{log}(p_{i}(\omega_{c})),$ (9)
where $\alpha,\beta$ are the hyperparameters that control the balance between the main loss and the auxiliary cross-entropy losses.
E  The model comparison results in metrics (specificity, recall, sensitivity, and NPV)
@@
BERT x 0.1767$\pm$0.0030 0.6966$\pm$0.0225 0.5634$\pm$0.0217 0.9330$\pm$0.0021
BioBERT x 0.1793$\pm$0.0011 0.6029$\pm$0.0278 0.6299$\pm$0.0194 0.9225$\pm$0.0028
Clinical Bert x 0.1962$\pm$0.0012 0.7158$\pm$0.0080 0.6071$\pm$0.0073 0.9410$\pm$0.0009
Multimodal (BERT) x x 0.2761$\pm$0.0034 0.7807$\pm$0.0063 0.7255$\pm$0.0064 0.9611$\pm$0.0008
Multimodal (BioBERT) x x 0.2741$\pm$0.0031 0.7934$\pm$0.0065 0.7182$\pm$0.0067 0.9629$\pm$0.0008
Multimodal (Clinical Bert) x x 0.2908$\pm$0.0027 0.7678$\pm$0.0063 0.7489$\pm$0.0054 0.9601$\pm$0.0008
Ours (MLP-BERT) x x 0.3297$\pm$0.0076 0.6957$\pm$0.0168 0.8094$\pm$0.0114 0.9522$\pm$0.0020
Ours (MLP-BioBERT) x x 0.3377$\pm$0.0115 0.6792$\pm$0.0218 0.8191$\pm$0.0148 0.9503$\pm$0.0023
Ours (MLP-Clinical Bert) x x 0.3609$\pm$0.0075 0.6478$\pm$0.0123 0.8454$\pm$0.0079 0.9472$\pm$0.0013
Ours (ResNet-BERT) x x 0.3179$\pm$0.0073 0.7206$\pm$0.0168 0.7916$\pm$0.0118 0.9550$\pm$0.0020
Ours (ResNet-BioBERT) x x 0.3244$\pm$0.0100 0.7092$\pm$0.0219 0.8000$\pm$0.0150 0.9538$\pm$0.0025
Ours (ResNet-Clinical Bert) x x 0.326$\pm$0.0067 0.7151$\pm$0.0131 0.8010$\pm$0.0097 0.9546$\pm$0.0015
Ours (FT-Trans-BERT) x x 0.3163$\pm$0.0080 0.7433$\pm$0.0109 0.7836$\pm$0.0110 0.9580$\pm$0.0012
Ours (FT-Trans-BioBERT) x x 0.3282$\pm$0.0015 0.7193$\pm$0.0055 0.8027$\pm$0.0027 0.9553$\pm$0.0007
Ours (FT-Trans-Clinical Bert) x x 0.3289$\pm$0.0070 0.7371$\pm$0.0171 0.7974$\pm$0.0108 0.9578$\pm$0.0021
(a) Mortality prediction
@@
BERT x 0.2004$\pm$0.0031 0.5685$\pm$0.0147 0.6488$\pm$0.0160 0.9070$\pm$0.0009
BioBERT x 0.2006$\pm$0.0022 0.5504$\pm$0.0160 0.6610$\pm$0.0141 0.9051$\pm$0.0013
Clinical Bert x 0.2208$\pm$0.0024 0.6251$\pm$0.0133 0.6587$\pm$0.0116 0.9193$\pm$0.0014
Multimodal (BERT) x x 0.2568$\pm$0.0029 0.6892$\pm$0.0141 0.6915$\pm$0.0108 0.9347$\pm$0.0022
Multimodal (BioBERT) x x 0.2554$\pm$0.0034 0.6943$\pm$0.0136 0.6868$\pm$0.0114 0.9358$\pm$0.0018
Multimodal (Clinical Bert) x x 0.2507$\pm$0.0045 0.7201$\pm$0.0188 0.6664$\pm$0.0163 0.9394$\pm$0.0024
Ours (MLP-BERT) x x 0.2809$\pm$0.0038 0.6218$\pm$0.0050 0.7538$\pm$0.0064 0.9281$\pm$0.0005
Ours (MLP-BioBERT) x x 0.2892$\pm$0.0048 0.6022$\pm$0.0061 0.7710$\pm$0.0071 0.9262$\pm$0.0007
Ours (MLP-Clinical Bert) x x 0.2980$\pm$0.0075 0.5899$\pm$0.0134 0.7841$\pm$0.0117 0.9254$\pm$0.0012
Ours (ResNet-BERT) x x 0.2787$\pm$0.0071 0.6460$\pm$0.0158 0.7402$\pm$0.0148 0.9314$\pm$0.0016
Ours (ResNet-BioBERT) x x 0.2848$\pm$0.0059 0.6212$\pm$0.0126 0.7580$\pm$0.0119 0.9285$\pm$0.0012
Ours (ResNet-Clinical Bert) x x 0.2968$\pm$0.0029 0.6227$\pm$0.0073 0.7720$\pm$0.0058 0.9299$\pm$0.0008
Ours (FT-Trans-BERT) x x 0.2595$\pm$0.0036 0.7180$\pm$0.0116 0.6830$\pm$0.0110 0.9402$\pm$0.0014
Ours (FT-Trans-BioBERT) x x 0.2668$\pm$0.0042 0.6914$\pm$0.0106 0.7059$\pm$0.0109 0.9368$\pm$0.0012
Ours (FT-Trans-Clinical Bert) x x 0.2782$\pm$0.0026 0.6766$\pm$0.0046 0.7288$\pm$0.0050 0.9359$\pm$0.0005
(b) PLOS prediction
Table 3: The model comparison for mortality (a) and PLOS prediction (b) tasks in metrics (precision, recall, specificity, and NPV).
F  The evaluation on different fusion settings for PLOS prediction
@@
Figure 7: The evaluation of our framework on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.

### v2

@@ 5.3 Evaluation on different fusion settings
Regarding predictive accuracy, the models achieved lower BACC across the three fusion settings for both tasks. Notably, the model based on modalities outperformed those based on data types and data sources in F1, AUROC, and AUPRC metrics. In terms of reliability evaluation metrics, the model based on modalities demonstrated superior reliability across three different structured data encoders. Additionally, Figure 4 and 5 suggest a positive relationship between predictability and reliability in our framework.
@@
Figure 4: The evaluation of our framework using Clinical BERT as text encoder on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
Figure 5: The evaluation of our framework using Clinical BERT as text encoder on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.
6 DISCUSSION
@@
In this section, we present our findings across four key areas: the advantages of using free-text notes for mortality and PLOS prediction, the effectiveness of the evidence-based multimodal framework, the impact of various fusion settings, and the influence of encoder selection.
6.1 Benefits of free-text notes
@@
From Tables 2 and 3, it is evident that integrating free-text notes with structured data enhances ICU outcome prediction, boosting predictability. Free-text notes provide information not included in structured EHR data, such as nursing details, physician documentation, and radiology reports after ICU admission. This suggests that free-text EHR notes and structured inputs can complement each other in predictive modeling, leading to improved performance.
6.2 Effectiveness on multimodal evidence fusion
Accurate and reliable ICU decision support
The proposed framework outperforms existing multimodal approaches by leveraging belief function theory for the effective fusion of structured and unstructured EHR data. This enables more accurate and robust predictions, which is essential for clinical decision-making in ICU settings. Notably, while the improvement in BACC is not significant (1.05% for mortality and 1.02% for PLOS prediction), the F1 score shows a significant increase (9.74% for mortality and 6.04% for PLOS prediction). This highlights the ability of the framework to identify critical ICU cases, as the F1 score prioritizes precision and recall, making it especially valuable in imbalanced datasets where positive cases are rare. Given the high-stakes nature of ICU outcomes, this improvement suggests our model enhances early identification of high-risk patients, potentially supporting timelier interventions in critical care.
Unlike traditional multimodal approaches, our evidence-based framework demonstrates greater prediction reliability, which is particularly valuable in clinical decision-making. By effectively handling uncertainty and inconsistencies in patient data under the proposed multimodal fusion framework, our approach ensures more trustworthy predictions with lower Brier and NLL (shown in Table 4 and 5). This enhanced reliability is crucial for ICU decision support, where inaccurate predictions can lead to overuse of critical resources or missed early interventions.
Efficient resource allocation
@@ 6.2 Effectiveness on multimodal evidence fusion
The experimental results in Tables 6 and 7 demonstrate that while existing multimodal approaches can effectively capture true positive instances, they often come at the cost of increased false positives. In ICU settings, such false positives can lead to the unnecessary use of medical resources and equipment. In contrast, our framework achieves a better balance by demonstrating higher precision and specificity, effectively reducing false positives. This capability is crucial for ensuring that ICU resources are allocated appropriately to patients in critical need.
@@
BERT as text encoder
Text encoder only x 0.1767$\pm$0.0030 0.6966$\pm$0.0225 0.5634$\pm$0.0217 0.9330$\pm$0.0021
Multimodal x x 0.2761$\pm$0.0034 0.7807$\pm$0.0063 0.7255$\pm$0.0064 0.9611$\pm$0.0008
[…]
Ours (ResNet) x x 0.3179$\pm$0.0073 0.7206$\pm$0.0168 0.7916$\pm$0.0118 0.9550$\pm$0.0020
Ours (FT-Trans) x x 0.3163$\pm$0.0080 0.7433$\pm$0.0109 0.7836$\pm$0.0110 0.9580$\pm$0.0012
BioBERT as text encoder
Text encoder only x 0.1793$\pm$0.0011 0.6029$\pm$0.0278 0.6299$\pm$0.0194 0.9225$\pm$0.0028
Multimodal x x 0.2741$\pm$0.0031 0.7934$\pm$0.0065 0.7182$\pm$0.0067 0.9629$\pm$0.0008
[…]
Ours (ResNet) x x 0.3244$\pm$0.0100 0.7092$\pm$0.0219 0.8000$\pm$0.0150 0.9538$\pm$0.0025
Ours (FT-Trans) x x 0.3282$\pm$0.0015 0.7193$\pm$0.0055 0.8027$\pm$0.0027 0.9553$\pm$0.0007
Clinical BERT as text encoder
Text encoder only x 0.1962$\pm$0.0012 0.7158$\pm$0.0080 0.6071$\pm$0.0073 0.9410$\pm$0.0009
Multimodal x x 0.2908$\pm$0.0027 0.7678$\pm$0.0063 0.7489$\pm$0.0054 0.9601$\pm$0.0008
[…]
Ours (ResNet) x x 0.3260$\pm$0.0067 0.7151$\pm$0.0131 0.8010$\pm$0.0097 0.9546$\pm$0.0015
Ours (FT-Trans) x x 0.3289$\pm$0.0070 0.7371$\pm$0.0171 0.7974$\pm$0.0108 0.9578$\pm$0.0021
Table 6: Comparison of class-specific prediction accuracy on mortality prediction.
@@
BERT as text encoder
Text encoder only x 0.2004$\pm$0.0031 0.5685$\pm$0.0147 0.6488$\pm$0.0160 0.9070$\pm$0.0009
Multimodal x x 0.2568$\pm$0.0029 0.6892$\pm$0.0141 0.6915$\pm$0.0108 0.9347$\pm$0.0022
[…]
Ours (ResNet) x x 0.2787$\pm$0.0071 0.6460$\pm$0.0158 0.7402$\pm$0.0148 0.9314$\pm$0.0016
Ours (FT-Trans) x x 0.2595$\pm$0.0036 0.7180$\pm$0.0116 0.6830$\pm$0.0110 0.9402$\pm$0.0014
BioBERT as text encoder
Text encoder only x 0.2006$\pm$0.0022 0.5504$\pm$0.0160 0.6610$\pm$0.0141 0.9051$\pm$0.0013
Multimodal x x 0.2554$\pm$0.0034 0.6943$\pm$0.0136 0.6868$\pm$0.0114 0.9358$\pm$0.0018
[…]
Ours (ResNet) x x 0.2848$\pm$0.0059 0.6212$\pm$0.0126 0.7580$\pm$0.0119 0.9285$\pm$0.0012
Ours (FT-Trans) x x 0.2668$\pm$0.0042 0.6914$\pm$0.0106 0.7059$\pm$0.0109 0.9368$\pm$0.0012
Clinical BERT as text encoder
Text encoder only x 0.2208$\pm$0.0024 0.6251$\pm$0.0133 0.6587$\pm$0.0116 0.9193$\pm$0.0014
Multimodal x x 0.2507$\pm$0.0045 0.7201$\pm$0.0188 0.6664$\pm$0.0163 0.9394$\pm$0.0024
[…]
Ours (ResNet) x x 0.2968$\pm$0.0029 0.6227$\pm$0.0073 0.7720$\pm$0.0058 0.9299$\pm$0.0008
Ours (FT-Trans) x x 0.2782$\pm$0.0026 0.6766$\pm$0.0046 0.7288$\pm$0.0050 0.9359$\pm$0.0005
Table 7: Comparison of class-specific prediction accuracy on PLOS prediction.
6.3 Analysis on different fusion settings
To explore the performance of our framework across different fusion settings, it is evident that models based on modalities achieved higher F1 scores but lower BACCs compared to the other two fusion settings. This discrepancy arises from the metrics’ focus: the F1 score emphasizes performance on the positive class by balancing precision and recall, while BACC provides an overall assessment of recall across all classes. Thus, models based on modalities are more effective at identifying ICU outcomes, likely due to the enhanced integration of information from independent sources facilitated by belief function theory. This observation aligns with our analysis of data independence. Figures 6 and 7 visualize the independence of structured features and four types of free-text notes, respectively. In Figure 6, correlation coefficients confirm that features within structured data are not independent. Meanwhile, Figure 7 shows that Radiology notes form a distinct cluster, while the other types of notes overlap, indicating a lack of independence among them. Moreover, the difference in BACC performance using the FT-Transformer across three fusion settings is smaller than with the MLP. This is likely due to the stronger predictive capability of the FT-Transformer.
6.4 Influence of encoders selection
The evaluations presented in Table 2 and 3 reveal that the choice of encoders plays a critical role in determining predictability. Among the assessed models, the FT-Transformer stands out as highly effective for structured tabular data, aligning with the findings of [34]. This effectiveness can be attributed to the transformer’s capability to capture complex relationships among transformed numerical and categorical features, which enhances its predictive power.
7 CONCLUSION
In this paper, we address the challenge of accurately and reliably predicting ICU outcomes by introducing a multimodal framework based on belief function theory that models both structured EHR data and free-text EHR notes. Our framework transforms deep features extracted from these two modalities into evidence through the evidence mapping module, which is then fused using Dempster’s rule to make final predictions. Through experiments on the MIMIC-III dataset, we demonstrate the effectiveness of our framework in terms of predictive accuracy and reliability. The study highlights its capability in managing heterogeneous multimodal EHR data, reducing false positives and potentially improving the allocation of medical resources in the ICU.
While this paper focuses on binary classification tasks, many clinical applications require solutions for multiclass tasks (e.g., disease diagnosis) and continuous regression tasks (e.g., survival prediction). These are equally important and relevant for advancing clinical practice. In the future, we plan to expand our framework by incorporating additional data modalities, such as time series and medical images, to provide deeper clinical insights and enhance model performance. We also aim to extend the framework to handle multimodal EHR multiclass tasks, offering valuable predictive guidance for complex clinical scenarios. Additionally, we intend to investigate regression tasks, leveraging the recently introduced Epistemic Random Fuzzy Set (ERFS) theory [46, 26] and further building on developments in evidential regression [24].
[…]
Acknowledgements
[…]
Code Availability The code implementation is available on Github repository (https://github.com/yuchengruan/evid_multimodal_ehr).
Declarations
[…]
Appendix A Evaluation on different fusion settings using BERT and BioBERT as text encoders
@@
Figure 8: The evaluation of our framework using BERT as text encoder on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
Figure 9: The evaluation of our framework using BERT as text encoder on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.
Figure 10: The evaluation of our framework using BioBERT as text encoder on different fusion settings for mortality prediction: (1) modalities, (2) data types, (3) data sources.
Figure 11: The evaluation of our framework using BioBERT as text encoder on different fusion settings for PLOS prediction: (1) modalities, (2) data types, (3) data sources.